Pollsters are concerned about declining levels of cooperation among persons contacted in surveys. A pollster contacts 73 people in the 18-21 age bracket and finds that 62 of them respond and 11 refuse to respond. When 305 people in the 22-29 age bracket are contacted, 281 respond and 24 refuse to respond. Assume that 1 of the 378 people is randomly selected. Find the probability of getting someone in the 22-29 age bracket or someone who agreed to respond.
Answer #1
respond refuse total
18-21 62 11 73
22-29 281 24 305
total 343 35 378

probability of getting someone in the 22-29 age bracket or someone who agreed to respond = P(22-29 age bracket) + P(respond) - P(respond in age bracket 22-29) = (305+343-281)/378 = 0.9709(answer)